OpenStudy (anonymous):

What are the solutions of x2 + 5x − 24 = 0? Show all work for full credit.

OpenStudy (anonymous):

do it by factorizing...

hartnn (hartnn):

hint : split the middle term as 8x-3x

OpenStudy (anonymous):

break 5x in two terms.

OpenStudy (anonymous):

8x-3x

OpenStudy (anonymous):

would it be −b±/b^2 - 4ac ------------- 2a

hartnn (hartnn):

thats using Quadratic formula...are u asked to use that ? if not factorizing will be better approach

OpenStudy (anonymous):

ya its asking for quadratic formula

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find a,b,c then the two roots of x are: \(\huge{x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

hartnn (hartnn):

so, first tell me what u get for a,b,c.

OpenStudy (anonymous):

a=x^2 b=5 c=24

hartnn (hartnn):

no, 'a' is the co-efficient of x^2 here u have 1x^2 so a=.... ?

OpenStudy (anonymous):

1^2

hartnn (hartnn):